OpenStudy (anonymous):

How many solutions are there to the equation y = 14x – 9? Explain your answer. A. There are no solutions because no value can be used for the variable x for the equation to remain true. B. There is one solution because only one value can be used for the variable x for the equation to remain true. C. There are 2 solutions because only 2 values can be used for the variables for the equation to remain true. D. There are an infinite number of solutions because any value can be used for the variable x and the equation would remain true.

OpenStudy (anonymous):

There are an infinite number of solutions because any value can be used for the variable x and the equation would remain true.

OpenStudy (anonymous):

@welshfella @mathmate Is D correct?And why?.

OpenStudy (welshfella):

yes D is correct we could be still putting possible solutions in next year if we wanted too

OpenStudy (anonymous):

I need help with 4 more!:)

OpenStudy (welshfella):

y = 14x - 9 put x = 1 then y = 14-9 = 5 x = 2 then y = 28-9 = 19 etc etc etc we could put any real number for x and find y

OpenStudy (anonymous):

OHHHH I see thanks :)

OpenStudy (welshfella):

OK , but could you plz put them in separate posts